### v1

# TMT: Tri-Modal Translation between Speech, Image, and Text by Processing Different Modalities as Different Languages

Abstract
The capability to jointly process multi-modal information is becoming an essential task. However, the limited number of paired multi-modal data and the large computational requirements in multi-modal learning hinder the development. We propose a novel Tri-Modal Translation (TMT) model that translates between arbitrary modalities spanning speech, image, and text. We introduce a novel viewpoint, where we interpret different modalities as different languages, and treat multi-modal translation as a well-established machine translation problem. To this end, we tokenize speech and image data into discrete tokens, which provide a unified interface across modalities and significantly decrease the computational cost. In the proposed TMT, a multi-modal encoder-decoder conducts the core translation, whereas modality-specific processing is conducted only within the tokenization and detokenization stages. We evaluate the proposed TMT on all six modality translation tasks. TMT outperforms single model counterparts consistently, demonstrating that unifying tasks is beneficial not only for practicality but also for performance.
1 Introduction
$*{}^{*}$qual contribution.
In search of artificial general intelligence, seamlessly processing and representing multi-modal information is a crucial prerequisite. Leveraging numerous efforts on the investigation of diverse multi-modal work, systems capable of processing audio-visual Chen et al. (2021); Hong et al. (2023), audio-text Yang et al. (2023); Wu et al. (2024), and text-visual Ramesh et al. (2021); Li et al. (2023b) modalities are actively emerging. Furthermore, developing a unified model that processes tri-modalities (i.e., audio, visual, and text) is drawing attention given its potential for advancing general intelligence Wu et al. (2023); Han et al. (2023).
However, building such a unified model remains challenging due to multifaceted factors. The distinct nature of different modalities intrinsically burdens the multi-modal systems to include two roles: processing modality-specifics and translating between different modalities. Consequently, the computational load required to train multi-modal models usually exceeds that of uni-modal models, attributed to the effort necessary to do two things simultaneously [^1] We believe this computational burden is one of the most impactful bottlenecks that hinder large multi-modal models, compared to Large Language Models (LLMs), where one can scale the model size and efficiently train it with text-only data.. Furthermore, the necessity of paired multi-modal data is another obstacle because the amount of such data (e.g., speech-image-text) is less abundant than uni-modal data.
In this paper, we propose a novel Tri-Modal Translation (TMT) framework that freely translates between different modalities (i.e., speech, image, and text) using a single model. Specifically, we present a novel viewpoint on multi-modal modeling by interpreting different modalities as different languages. We split the Multi-Modal Translation (MMT) problem into modality-specific processings and an integrated translation, where the latter is treated as the Neural Machine Translation (NMT). To this end, we discretize all modalities into tokens in advance [^2] Image and speech are represented as sequences of integers. by employing pre-trained modality-specific tokenizers. This enables us to leverage the expertise from each field, discarding modality-specifics from the core translation modeling.
The proposed TMT framework and viewpoint introduce several benefits. First, the proposed system is “purely multi-modal” by uniformly interfacing all modalities together different to the method of mapping different modalities into the text space of LLM Wu et al. (2023), which we believe is the future direction of building multi-modal LLMs. Second, through discretization, the computational burden in modeling speech and image is significantly relaxed, specifically by more than $99$Chang et al. (2023b); Kim et al. (2024a) in the aspect of required bits. Finally, we can employ unlimited training data through Back Translation (BT) Sennrich et al. (2016), a well-known data augmentation strategy in NMT333 BT typically translates target monolingual data into a source language to construct source-to-target parallel data.. Therefore, even utilizing uni-modal data Deng et al. (2009); Kahn et al. (2020); Fan et al. (2019) in building TMT becomes possible. With these advantages, we can efficiently scale up multi-modal training, combining computational efficiency with large-scale data.
The contributions of this paper can be summarized as follows: 1) To the best of our knowledge, this work is the first to explore translations between tri-modalities, image, speech, and text, by discretizing all modalities. 2) We provide a novel interpretation of modalities as languages. By regarding MMT as NMT, we show that target uni-modal corpora can be also employed through back translation. 3) The proposed TMT encompasses direct speech-to-image synthesis and image-to-speech captioning, both of which have not well been addressed in the previous literature. 4) We show that the proposed TMT can be efficiently trained like text-only systems via extensive experiments on six MMT tasks. Project page is available on github.com/ms-dot-k/TMT.
Figure 1: Illustration of the overall proposed Tri-modal Translation (TMT) model. (a) Image, speech, and text modalities are tokenized into discrete tokens, respectively, using modality-specific pre-trained tokenizers. (b) The TMT model encodes any of the three modalities and translates the encoded representation into a desired output modality. All modalities are represented in discrete tokens and hence our “multi-modal” translation (MMT) task is treated as a neural machine translation (NMT) task between different “languages”, with improved efficiency.
2 Related Work
NMT aims to transform a source language sentence into a target language Stahlberg (2020). With advanced network architectures Kenton and Toutanova (2019); Lewis et al. (2020) and training methodologies, including denoising pre-training Ho et al. (2020) and BT Sennrich et al. (2016), the field has rapidly developed. In this work, we interpret discretized speech and images as different languages so that TMT can be addressed using the format of NMT. In this viewpoint, NMT techniques, back translation and network architecture, are incorporated into TMT.
MMT. Developing models capable of performing multiple MMT tasks is a popular research topic with diverse groups of researchers’ investigation Kim et al. (2023); Ao et al. (2022); Girdhar et al. (2023). MUGEN Hayes et al. (2022) explored the discretization of video, audio, and text obtained from synthetic game data, and successfully trained multiple MMT models capable of a single MMT task. VoxtLM Maiti et al. (2024) proposed a decoder-only architecture to train a model that performs four MMT tasks ($2\times 2$) spanning speech and text modalities. They discretized speech using a pre-trained self-supervised feature and treated discretized speech tokens in the same manner as text tokens within a combined vocabulary set. SEED Ge et al. (2023a, b) proposed an image tokenizer using contrastive learning between image and text. With the learned image tokens, they finetuned an LLM spanning text and image modalities, covering two MMT tasks, image captioning and text-to-image synthesis. Different to VoxtLM and SEED, TMT is not restricted to bi-modal translation and can translate between tri-modalities.
There also exist a few concurrent works aiming to handle tri-modalities. NExT-GPT Wu et al. (2023) tries to adapt different modalities into the text space of a pre-trained LLM by using several modality-specific encoders and adaptors, where the overall performance depends on the performance of the LLM. Distinct from NExT-GPT, our TMT is “born multi-modal” by learning from multi-modal data instead of projecting different modalities’ representation into that of a text LLM. Gemini Team et al. (2023) is a recently proposed multi-modal LLM. Different to Gemini, the proposed TMT can directly synthesize speech. Most significantly, different to the existing work, TMT treats different modalities as different languages by interpreting discretized image and speech data equivalent to text tokens. Hence, modality specifics are separated and all six MMT tasks are jointly modeled with a unified interface with fully shared parameters.
3 Method
[…]
3.1 Individual tokenization of each modality
Discretized representations have demonstrated promising results in various research fields. In image processing, Vector Quantization (VQ) Van Den Oord et al. (2017); Esser et al. (2021); Ge et al. (2023a, b) has successfully expressed images using discrete tokens. In speech processing, VQ Défossez et al. (2022); Wang et al. (2023) and Self-Supervised Learning (SSL) Baevski et al. (2020); Hsu et al. (2021a); Chen et al. (2022) model-based discrete tokens Lakhotia et al. (2021) are showing remarkable performances across different tasks (e.g., speech translation Lee et al. (2022b); Popuri et al. (2022); Choi et al. (2023) and synthesis Polyak et al. (2021); Hayashi and Watanabe (2020)).
Motivated by the recent success, we tokenize image, speech, and text modalities. Specifically, as shown in Fig. 1a, the raw image is tokenized into image tokens $u_{i}\in\mathbb{N}^{L_{i}}$ by using SEED-2 Ge et al. (2023b) tokenizer, where each image is compressed into 32 tokens (i.e., $L_{i}=32$). Speech is also tokenized into discrete values $u_{s}\in\mathbb{N}^{L_{s}}$ by clustering the features of a HuBERT Hsu et al. (2021a), an SSL speech model, whose output is 50Hz. Following the practices in Lakhotia et al. (2021), we merge duplicated tokens so that the granularity of speech tokens (i.e., $L_{s}$) is less than 50Hz. For the text input, we employ the BERT tokenizer Schuster and Nakajima (2012); Kenton and Toutanova (2019) to derive text tokens $u_{t}\in\mathbb{N}^{L_{t}}$.
Advantages of tokenizing tri-modalities. Tokenizing all three modalities brings two main advantages. Firstly, the MMT problem now has a unified interface. Modality specifics are processed mostly within tokenization and detokenization process.
  Modality Data Format Data Size (bits) Data Size (%)
Speech Raw Audio 16000 $\times$ L $\times$ 16 100%
Mel-spectrogram 100 $\times$ 80 $\times$ L $\times$ 32 100%
Speech Token ($<$50) $\times$ L $\times$ 8 $<$0.2%
Image Raw Image 224 $\times$ 224 $\times$ 3 $\times$ 8 100%
Image Token 32 $\times$ 13 0.035%
Text Raw Text S $\times$ 8 -
Table 1: Data size (bits) comparisons between raw data and tokens. The values are based on the 16kHz 16bits for raw speech waveforms, 224$\times$224 for raw RGB image, 80 filterbanks with 10ms shift for mel-spectrogram. $L$ is the length (sec) of the audio. For text, one character is assigned 8 bits. $S$ refers to the length of the characters.
The other key advantage is efficiency Chang et al. (2023b); Park et al. (2023). Processing raw image and audio is burdensome due to the necessity of large-scale data storage and computational resources Kim et al. (2024b). Discretizing speech and image reduces the bits required to less than 0.2% and 0.04% as shown in Table 1 and appendix F. Therefore, by tokenizing all modalities before training, we can significantly increase the computational efficiency for both storage and computing memory.
3.2 Integrated modeling of tri-modalities
As shown in Fig. 1b, we leverage a multi-modal encoder-decoder architecture Vaswani et al. (2017); Lewis et al. (2020); Liu et al. (2020), widely used in the NMT literature as the backbone architecture. Therefore, all parameters are shared between all six MMT tasks. A popular auto-regressive training scheme is employed. To inform both the encoder and the decoder about input and output modalities, we add a modal-type embedding to the embedded features after the token embedding layer, similar to language embeddings Conneau and Lample (2019). Then, the losses for each MMT are summed and backpropagated. Formally put, the objective function of TMT training can be expressed as:
$\mathcal{L}=-\sum_{\begin{subarray}{c}m\in\mathcal{M}\\ m\neq k\end{subarray}}\ \sum_{k\in\mathcal{M}}\ \sum^{L_{m}}_{l=1}\ \log p(u_{m}^{l}|u_{m}^{<l},u_{k}),$ (1)
where $k$ and $m$ refer to a pair of non-overlapping modalities, $\mathcal{M}\,$=$\,\{i,s,t\}$ is the set of tri-modalities, and $u^{<l}$ represents the predictions before $l$-th frame.
3.3 MMT tasks covered by TMT
TMT incorporates all six MMT tasks (${}_{3}\mathbf{P}_{2}$) spanning different combinations of input and output modalities. Related works for each MMT task can be found in appendix E.
Image Captioning Xu et al. (2015) is a task of describing input image content using natural language. A comprehensive understanding of the input image and the naturalness of the output text are important.
Figure 2: Sample illustrations of the six tasks that the proposed TMT is capable of. (a) Image captioning, (b) Image-to-speech captioning, (c) Text-to-image synthesis, (d) Speech-to-image synthesis, (e) Automatics speech recognition, and (f) Text-to-speech synthesis. Note that we also show transcriptions, obtained using a pre-trained ASR system, of speech to ease understanding.
  Task Methods COCO Flickr8k
BLEU-4 METEOR ROUGE CIDEr SPICE BLEU-4 METEOR ROUGE CIDEr SPICE
Image normal-$\rightarrow$ Text Single MMT 28.9 27.3 53.6 101.9 20.5 21.8 24.3 49.7 67.7 19.1
TMT 31.4 28.1 54.9 108.7 21.3 26.0 25.8 52.5 79.7 20.5
Image normal-$\rightarrow$ Speech Single MMT 22.5 21.5 47.1 69.1 15.0 16.9 20.0 45.6 44.7 13.5
TMT 24.7 23.2 48.9 78.7 16.7 21.3 22.2 48.5 55.2 15.4
Table 2: Comparisons between single MMT systems and TMT on the image and speech-driven captioning tasks. All models are trained on combined COCO and Flickr8k train sets.
  Task Methods COCO Flickr8k
CLIP WER CLIP WER
Text normal-$\rightarrow$ Image Single MMT 68.8 - 69.3 -
TMT 68.2 - 68.1 -
Speech normal-$\rightarrow$ Image Single MMT 57.5 - 58.5 -
TMT 67.3 - 67.0 -
Speech normal-$\rightarrow$ Text Single MMT - 6.9 - 6.8
TMT - 6.5 - 6.1
Text normal-$\rightarrow$ Speech Single MMT - 16.0 - 9.7
TMT - 11.8 - 9.5
Table 3: Comparisons between single MMT systems and TMT on four MMT tasks. All models are trained on combined COCO and Flickr8k train sets.
Image-to-Speech Captioning Hsu et al. (2021b) is a task that directly synthesizes a spoken description of the input image. This task is used for the image-to-speech translation evaluation of TMT. It is regarded as a challenging task due to insufficient supervision and the complex nature of speech.
Text-driven Image Synthesis Ramesh et al. (2021, 2022); Saharia et al. (2022); Chang et al. (2023a) aims to synthesize an image that matches a given text description. Both the naturalness of the generated image and the correctness of description are important.
Speech-driven Image Synthesis generates an image from a spoken description Li et al. (2020); Wang et al. (2020). The task is equal to solving an automatic speech recognition (ASR) and text-driven image synthesis at once without the intermediate text representation. Similar to the relationship between image captioning and image-to-speech captioning, further challenges exist, mainly due to the absence of text modality. This task remains crucial for those languages that do not have writing systems Lee et al. (2022b), where text-based systems cannot be applied. Note that this paper is the first to synthesize high-resolution images of 768 $\times$ 768 from speech.
ASR is a well-established task that generates text transcriptions from an input speech Amodei et al. (2016); Prabhavalkar et al. (2023). We use this task for speech-to-text translation of TMT.
Text-to-Speech Synthesis (TTS) generates spoken content corresponding to the given text transcription Wang et al. (2017); Tan (2023).
3.4 Back Translation (BT)
[…]
4 Experiments
4.1 Datasets
The training data comprises Conceptual Captions 3M (CC3M), Conceptual Captions 12M (CC12M) Sharma et al. (2018); Changpinyo et al. (2021), COCO Lin et al. (2014), SpokenCOCO Hsu et al. (2021b), Flickr8k Hodosh et al. (2013), and Flickr8kAudio Harwath and Glass (2015). For COCO and Flickr8k, we employ the original corpora for image-text pairs and then further employ SpokenCOCO and Flickr8kAudio, their recorded speech version, to comprise audio-text-image tri-modal pairs. For CC3M and CC12M, VITS Kim et al. (2021a), a TTS model trained on VCTK Yamagishi et al. (2019), is employed to synthesize speech from random speakers to compile audio-text-image pairs. The evaluation is performed on the test split of COCO and Flickr8k after finetuning. The popular Karpathy splits Karpathy and Fei-Fei (2015) are employed for COCO and Flickr8k.
4.2 Metrics
We leverage diverse metrics to evaluate six MMT tasks spanning three output modalities.
Captioning tasks. For the image captioning and image-to-speech captioning tasks, we employ BLEU Papineni et al. (2002), METEOR Denkowski and Lavie (2014), ROUGE Lin (2004), CIDEr Vedantam et al. (2015), and SPICE Anderson et al. (2016). BLEU, METEOR, ROUGE, and CIDEr measure n-gram overlaps between predictions and references. SPICE compares semantic propositional content. Across all metrics, higher value indicates better performance.
Image synthesis tasks. For the text- and speech-driven image synthesis tasks, we adopt the CLIP score Radford et al. (2021) (i.e., the cosine similarity) to assess how similar the generated image is associated with the input text or speech, following Ge et al. (2023a); Koh et al. (2023).
ASR and TTS. We employ the widely used Word Error Rate (WER) for ASR and TTS. For TTS, we transcribe the generated speech by using a pre-trained ASR model Baevski et al. (2020), following Lakhotia et al. (2021); Kim et al. (2023). Additionally, we employ Mean Opinion Score (MOS) and neural MOS to assess the quality of generated results.
  Input Representation Methods Num. MMT. COCO Flickr8k
BLEU-4 METEOR ROUGE CIDEr SPICE BLEU-4 METEOR ROUGE CIDEr SPICE
Raw Image SAT Xu et al. (2015) 1 24.3 23.9 - - - 21.3 20.3 - - -
BLIP-2 Li et al. (2023a) 1 43.7 - - 145.8 - - - - - -
Discrete Image Token Kim et al. (2024a) 1 29.9 25.2 52.8 97.4 18.6 23.4 22.0 48.9 63.3 15.4
Ge et al. (2023a) 2 34.6 28.4 56.4 131.0 22.9 - - - - -
[…]
Table 4: Image captioning performance (Image-to-Text Translation) comparisons on COCO and Flickr8k. We also report the performance of raw image-based models for reference purposes. “Num. MMT.” indicates the number of MMT tasks that the model is capable of.
  Input Representation Methods Num. MMT. COCO Flickr8k
BLEU-4 METEOR ROUGE CIDEr SPICE BLEU-4 METEOR ROUGE CIDEr SPICE
Raw Image Wang et al. (2021) 1 - - - - - 3.5 11.3 23.2 8.0 -
Hsu et al. (2021b) 1 23.3 21.2 47.8 73.2 14.9 12.5 14.5 39.1 24.5 9.5
Effendi et al. (2021) 1 - - - - - 14.8 17.4 32.9 45.8 -
Kim et al. (2024a) 1 25.9 23.8 50.4 81.1 17.5 20.6 22.0 48.4 53.6 15.8
Discrete Image Token Kim et al. (2024a) 1 20.1 21.4 46.4 64.0 15.0 16.7 19.6 44.2 41.2 13.1
[…]
Table 5: Image-to-speech captioning performance (Image-to-Speech Translation) comparisons on COCO and Flickr8k. We also report the performance of raw image-based models for reference purposes.
4.3 Implementation Details
The input image is resized into 224 $\times$ 224, while the output image size is 768 $\times$ 768. The input audio is resampled to 16kHz and tokenized using a pre-trained HuBERT base model Hsu et al. (2021a). The backbone model that models the multi-modal information is an encoder-decoder Transformer Vaswani et al. (2017) with 12 layers each. For pre-training the TMT before applying BT, we train the model for 350k steps. Then, we further train the model using the additional data obtained by applying BT, for 250k steps with the same configuration but with a peak learning rate of $5e^{-5}$. For decoding, we use beam search with a beam width of 5. Full implementation details including hyper-parameters can be found in appendix B.
4.4 Results
Firstly, we illustrate selected samples for each task of TMT in Fig. 2. We then compare single MMT task models versus the proposed TMT model. Next, we compare our TMT with other works in the literature for each MMT task. Finally, we conduct a human subjective evaluation on two speech generation tasks: Image-to-Speech synthesis and TTS. Further ablations on the discrete token vocabulary size and image tokenizer are presented in appendix C.
4.4.1 Effectiveness of Unifying Tasks
We start with the question: “Can training a single model, capable of multiple MMT tasks, have synergy in performance on top of efficiency compared to training multiple single MMT models?” We train six single MMT models and one TMT model using the train sets of COCO and Flickr8k. For fairness, all models are trained for 300k iterations.
Table 2 compares the captioning performances, including image-to-text and image-to-speech translations. The results show that by unifying tasks, we can consistently obtain performance gain. In our analysis, this stems from the unified multi-modal decoder. The decoder can leverage complementary multi-modal information in learning language modeling. Table 3 shows the comparison results for the remaining four tasks. Similar to captioning, the unified TMT model outperforms its single-task counterparts consistently, except for the text-to-image translation. However, the performance differences between single MMT and TMT in text-to-image translation are marginal, while the performance gain obtained by unifying tasks is substantial for speech-to-image translation. We conclude that unifying tasks is beneficial not only for practicality by building a single model but also in the aspect of performance.
4.4.2 Comparison with the literature
We compare TMT’s performance on each MMT task with the other works in the literature. The TMT model in this section employs 10M training samples.
Image Captioning Tasks. Table 4 shows the results of image-to-text translation of TMT and other competing systems. We also report the performances of raw image-based systems for reference purposes. We observe that our TMT accomplishes comparable performances among state-of-the-art discrete image-based models. Notably, the proposed TMT is competitive with the LLM-based method Ge et al. (2023a) with a much smaller number of parameters, on COCO. Note that Ge et al. (2023a) is built on OPT 2.7B model Zhang et al. (2022) while the proposed TMT consists of a 270M multi-modal model. This shows the effectiveness of the proposed TMT compared to the method of finetuning LLM. Moreover, TMT can perform six MMT tasks while the previous methods can only perform two MMT tasks at maximum.
Table 5 shows the direct image-to-speech captioning performances. TMT outperforms the recently proposed state-of-the-art Kim et al. (2024a) on both COCO and Flickr8k databases and even achieves comparable performances with raw-image based method. Since the decoder of the proposed TMT is trained using multi-tasks with text simultaneously, it better models linguistics compared to the transfer learning from the text of Kim et al. (2024a).
Image Synthesis Tasks. The upper part of Table 6 displays the performances of text- and speech-driven image synthesis tasks. Following the literature, we evaluate the model on Flickr30k instead of using Flickr8k for the text-to-image synthesis tasks. The proposed TMT achieves the best CLIP similarity on both databases in text-to-image translation, indicating that TMT generates images that best describe the given text sentences. Note that the proposed TMT has much fewer model parameters (i.e., 270M) compared to Koh et al. (2023) and Ge et al. (2023a) which are based on LLMs, OPT 6.7B, and OPT 2.7B, respectively. TMT still outperforms the competitors with more diverse MMT task abilities. Also, the speech-to-image translation performance of TMT on COCO is 68.88, closer to that of text-to-image performance. We hence confirm that even if using speech directly as input, TMT correctly generates images described in the speech. We underline that this work is the first to explore high-resolution image synthesis at a resolution of 768$\times$768 from speech. In addition to the samples shown in Fig 2, appendix A.
ASR and TTS Tasks. The lower part of Table 6 shows the performances of ASR and TTS. For the evaluation purpose, we compare the ASR performance with a pre-trained Wav2Vec2.0 (W2V2) Baevski et al. (2020) model. The results show that TMT can correctly recognize the input speech into text by achieving lower WER compared to W2V2. For TTS, as the proposed model is a discrete token-based speech synthesis model, we compare the performance with the resynthesized speech (ReSyn) from the ground-truth speech tokens following Lakhotia et al. (2021); Kim et al. (2023). Results demonstrate that the TMT can generate the correct speech from input text without suffering in synthesizing background noises which inherently exist in the ground-truth speech tokens.
  Methods Num. MMT. Image Synthesis (CLIP)
COCO Flickr30k
$\bullet$ Text-driven Image Synthesis
Rombach et al. (2022) 1
68.43
65.40
Koh et al. (2023) 2
67.45
65.16
Ge et al. (2023a) 2
68.23
65.23
[…]
$\bullet$ Speech-driven Image Synthesis
TMT 6 68.88
  Methods Num. MMT. Speech Processing (WER)
COCO Flickr8k
$\bullet$ Automatic Speech Recognition (ASR)
W2V2 Baevski et al. (2020) 1
9.70
7.26
[…]
$\bullet$ Text-to-Speech Synthesis (TTS)
ReSyn Lakhotia et al. (2021) -
29.32
22.21
[…]
Table 6: Text- and Speech-driven image synthesis, ASR, and TTS (Text-to-Image, Speech-to-Image, Speech-to-Text, and Text-to-Speech Translation) performance comparisons on COCO and Flickr8k. For TTS, we report the WER by feeding the synthesized audio into a pre-trained ASR model following Kim et al. (2023).
4.4.3 Mean Opinion Score Comparisons
In order to assess the quality of generated speech in both image-to-speech captioning and text-to-speech tasks, we conduct human subjective study and neural MOS test using MOSNet Lo et al. (2018) and SpeechLMScore Maiti et al. (2023). Specifically, we enlisted 21 participants to evaluate the naturalness of the generated speech for both tasks, and for the image-to-speech captioning task, we also assessed descriptiveness, gauging the extent to which the speech accurately describes the input images. Following the previous works Hsu et al. (2021b); Kim et al. (2024a), 20 samples are used for the image-to-speech captioning task. For TTS, 20 samples and 100 samples are employed for human subjective study and neural MOS, respectively. The results are shown in Table 7. TMT achieves the best score for both human study and neural MOS scores in image-to-speech captioning. Moreover, we achieve comparable performances with VITS on TTS task. Through MOS evaluation, we can confirm the effectiveness of TMT in generating intelligible and natural speech.
  Methods MOS Neural MOS
Natural Descriptive MOSNet $\uparrow$ SpeechLM $\downarrow$
$\bullet$ Image-to-Speech Captioning
Hsu et al. (2021b) 2.70±${}_{\pm 0.38}$ 2.96±${}_{\pm 0.22}$ 2.51 4.23
Kim et al. (2024a) 3.71±${}_{\pm 0.25}$ 3.41±${}_{\pm 0.23}$ 4.30 4.16
[…]
$\bullet$ Text-to-Speech Synthesis (TTS)
VITS Kim et al. (2021a) 3.85±${}_{\pm 0.25}$ 4.64 4.02
[…]
Table 7: Human evaluation score (MOS) comparisons with 95% confidence interval, and Neural MOS scores in image-to-speech captioning and TTS on COCO.
  Methods Image Captioningnormal-$\uparrow$ Image-to-Speechnormal-$\uparrow$ ASRnormal-$\downarrow$
TMT w/o BT 56.7 48.9 5.24
TMT 56.9 49.5 5.15
  Text-to-Imagenormal-$\uparrow$ Speech-to-Imagenormal-$\uparrow$ TTSnormal-$\downarrow$
TMT w/o BT 69.48 68.67 9.78
TMT 69.57 68.88 9.54
Table 8: Effectiveness of training with additional data derived with Back Translation (BT) on COCO. ROUGE is reported for Image Captioning and Image-to-Speech Captioning tasks.
4.4.4 Effectiveness of Back Translation
Table 8 shows the result of training TMT with and without BT. By employing additional 4M back translated data, we observe no significant yet consistent improvement across tasks of TMT. This shows the potential for further improvement, since the amount of additional data BT can produce is limitless. In appendix D, we show the examples of back translated data on both uni-modal data and bi-modal data.
5 Conclusion
We introduced TMT, a novel tri-modal translation model between speech, image, and text modalities. We interpreted different modalities as new languages and approached the MMT task as an NMT task after tokenizing all modalities. Our experiments demonstrated that the three modalities could be successfully translated using a multi-modal encoder-decoder architecture, incorporating six MMT tasks in a single model. TMT outperformed single MMT model counterparts. Remarkably, TMT achieves comparable performances with LLM-based methods consisting of more than 2.7B parameters, whereas TMT only has 270M parameters.
Appendix A Additional Qualitative Results
Fig. 3 illustrates additional generated images by using the text and speech as inputs (i.e.Text-driven Image Synthesis and Speech-driven Image Synthesis).
Appendix B Implementation details
[…]
Appendix C Ablation on Speech and Image Tokens
To confirm the effectiveness of the different multi-modal tokens, we perform ablation studies by differing configurations for speech and image tokens. First, we differentiate the number of k-means clustering of the speech tokenizer to obtain different token sizes. Then, we train single MMT models of image-to-speech captioning and ASR using 200, 500, and 1000 token sizes. Table 9 shows the image-to-speech captioning and ASR performances. The results show that there is a trade-off in performances of ASR (WER) and image-to-speech captioning (BLEU, METEOR, ROUGE, CIDEr, SPICE). We confirm a consistent tendency where image-to-speech captioning performs better with a smaller number of speech tokens. In contrast, for ASR, employing a large number of tokens is beneficial. Since the performance gaps in ASR are smaller than those of image-to-speech captioning, we employ 200 speech tokens for all experiments.
Second, we investigate the different types of visual tokens among VQ-GAN-based Esser et al. (2021); Yu et al. (2021) and SEED-based Ge et al. (2023a, b). To check the effect of visual tokens only, we employ the same architecture and training configuration of Kim et al. (2024a) and only change the visual tokens. The performances of image captioning and image-to-speech captioning on COCO are shown in Table 10. In both tasks, SEED-based token achieves better performances. As SEED tokenizer is trained along with image-text association learning, its token achieves better performance in image-language association with fewer tokens (i.e., 784 vs. 32). Therefore, we employ SEED-based tokenizer for our experiments.
  # Token BLEU-4 METEOR ROUGE CIDEr SPICE WER
200 Tokens 22.5 21.5 47.1 69.1 15.0 6.9
500 Tokens 17.7 19.0 44.3 52.7 12.3 6.3
1000 Tokens 15.1 17.7 41.7 48.3 11.9 6.3
Table 9: Ablation study to confirm the effects of vocabulary size of speech tokens on Image-to-Speech Captioning (BLEU, METEOR, ROUGE, CIDEr, SPICE) and Automatic Speech Recognition (WER).
Table 10: Captioning tasks’ performance comparisons using different visual token types.
  Visual Token Type Image Captioning
BLEU-4 METEOR ROUGE CIDEr SPICE
VQ-GAN-based 29.9 25.2 52.8 97.4 18.6
SEED-based 33.0 27.1 55.5 117.2 20.7
  Visual Token Type Image-to-Speech Captioning
BLEU-4 METEOR ROUGE CIDEr SPICE
VQ-GAN-based 20.1 21.4 46.4 64.0 15.0
SEED-based 20.6 21.6 47.4 65.1 15.7
Appendix D Example of Back Translation
We display the data obtained via BT on ImageNet and CommonVoice. For visualization purposes, we also plot the mel-spectrogram of audio. Fig. 4 and Fig. 5 show the results of BT on ImageNet dataset. By employing image-only uni-modal dataset, we can augment the paired data of text-image and speech-image using BT. The back translated dataset is utilized to further train our intermediate TMT model on Text- and Speech-driven Image Synthesis (text-to-image and speech-to-image MMT) tasks. Fig. 6 and Fig. 7 show the results on CommonVoice dataset. With the text-audio bi-modal dataset, we can build tri-modal paired dataset. The back translated dataset is used for Image Captioning and Image-to-Speech Captioning (image-to-speech and image-to-text MMT tasks) along with ASR and TTS (text-to-speech and speech-to-text MMT tasks).
Figure 4: Examples of back translated data on ImageNet dataset. The source text data is generated from the target image of ImageNet, building text-to-image pairs.
Figure 5: Examples of back translated data on ImageNet dataset. The source speech data is generated from the target image of ImageNet, building speech-to-image pairs.
Figure 6: Examples of back translated data on CommonVoice dataset. The source image data is generated from the target text of CommonVoice, building image-to-text pairs.
Figure 7: Examples of back translated data on CommonVoice dataset. The source image data is generated from the target speech of CommonVoice, building image-to-speech pairs.
Appendix E Related Work for Each MMT
E.1 Image Captioning
Image Captioning Xu et al. (2015) is a task to describe input image content in natural language. It has shown remarkable progress with the development of computer vision and natural language processing technologies. Recent methods showed that vision-language pre-training can significantly improve image captioning performances over training the model from scratch Wang et al. (2022). For example, Radford et al. (2021); Jia et al. (2021) proposed cross-modal contrastive learning, and Li et al. (2021); Kim et al. (2021b); Li et al. (2022, 2023a) proposed multi-task pre-training which contains such as image-text matching, masked language modeling, and contrastive learning. Basically, the task can be regarded as an MMT problem of image-to-text translation.
Similar yet distinct from image captioning, image-to-speech captioning Hsu et al. (2021b) aims to directly generate a spoken description of the input image. As the task does not go through with text modality, it is regarded as a challenging problem due to insufficient supervision and the complex nature of speech. Recent methods tried to mitigate the challenge by representing speech in pseudo-text format and showed promising results Hsu et al. (2021b); Effendi et al. (2021); Kim et al. (2024a). This task falls in text-to-speech translation.
E.2 Language-driven Image Synthesis
With the great development of visual codebook representation Van Den Oord et al. (2017); Esser et al. (2021) and diffusion Ho et al. (2020); Rombach et al. (2022), image generation has become more and more realistic. Especially, Ramesh et al. (2021, 2022); Saharia et al. (2022); Chang et al. (2023a) showed that generative models can now freely create images based on text descriptions by conditioning the text representations into the decoding process, such as diffusion process. They are the representatives of text-to-image translation tasks.
Different from text-based image synthesis, direct speech-to-image synthesis Li et al. (2020); Wang et al. (2020) has not well been addressed. Since many languages have no writing systems Lee et al. (2022b), text-based systems cannot be applied to such languages. Hence, the demand for technology enabling spoken language-driven image synthesis, the speech-to-image translation, is expected to increase.
Appendix F Details for Data Size (bits) Comparison
Table 1 shows the data size comparisons in bits between different data types of each modality. By employing speech tokens rather than directly using raw audio or its Mel-spectrogram, we significantly reduce the data size to less than 0.2% compared to other formats. Consequently, 1-second speech tokens contain a similar number of bits as text comprising approximately 50 characters. Furthermore, by utilizing image tokens instead of raw images, only 0.035% of the bits are needed to represent the original image content. This substantial reduction equates one image to the text length of 52 characters. Therefore, by representing and employing the tokenized data instead of directly utilizing their raw format, we can efficiently train our TMT system like training a text-only system with reduced data storage requirements and computing resources.
Appendix G Full Results of Back Translation
  Task Method BLEU-4 METEOR ROUGE CIDEr SPICE
Im. normal-$\rightarrow$ Txt. TMT w/o BT 33.9 29.3 56.7 116.7 22.3
TMT 34.3 29.2 56.9 117.7 22.3
Im. normal-$\rightarrow$ Sp. TMT w/o BT 25.0 22.8 48.9 79.5 16.6
TMT 25.8 23.3 49.5 81.3 16.8
Table 11: Performance comparisons of with BT and without BT on Image captioning and Image-to-speech captioning performances on COCO.
Table 11 shows the performances of image captioning and image-to-speech captioning on COCO in all 5 metrics when the BT is applied and not. By employing pseudo multi-modal pair datasets obtained through BT, we can consistently improve the performances overall. Please note that we can produce an unlimited number of pseudo-pair data through the multi-modal BT which may improve the performance more.